OpenStudy (anonymous):

Change 2x + 5y = -30 to slope-intercept form and state the slope and y-intercept. You must show all work

OpenStudy (anonymous):

solve for y in terms of x and the constant y =-2/5x -6

OpenStudy (anonymous):

2x + 5y = -30 5y = -2x-30 divide all by 5 y=-2/5x - 6